Calculate the amount of charge represented by six million protonsCalculate the amount of charge represented by six million protons
1
Expert's answer
2020-09-28T08:11:26-0400

Charge of one proton, "q = 1.6 \times 10^{-19} C"

Charge on six million protons, "Q = 6 \times 10^6 \times 1.6 \times10^{-19} C = 9.6 \times 10^{-13} C"